OpenStudy (anonymous):

A box holds 42 pieces of red construction paper and 18 pieces of blue construction paper. A piece of paper is pulled out of the box at random and then placed back in the box. How many times can you expect that a piece of red paper will be pulled out of the box in 110 tries? A. 48 B. 55 C. 60 D. 77

OpenStudy (myth1111):

42 + 18 = 60 42/60 = 70% 0.7*110 = ___

OpenStudy (anonymous):

77

OpenStudy (myth1111):

Looks good.

OpenStudy (anonymous):

Thanks! I only have one more hope I pass!